# Metallicities of M Dwarf Planet Host Stars from Kepler, K2, and TESS observed by APOGEE: Trends with Exoplanetary Radii and Orbital Periods

### v2

@@ Abstract
One important property in studying the exoplanet population is the host star metallicity ([M/H]). In this study, we derived stellar metallicities and oxygen abundances for 45 M dwarf stars (with 65 exoplanets) using the near-infrared high-resolution spectra from the SDSS APOGEE survey and synthetic spectra computed in LTE. We also investigated the exoplanetary radii distribution for a larger sample of 246 exoplanets orbiting 188 M dwarf stars. The [M/H] versus [O/M] distribution obtained indicates that our sample is composed mainly of thin disk stars, which follow the behavior of the low-alpha sequence in the Milky Way thin disk. Small planets with radii smaller than 3R⊕ were found around stars with a range of metallicities (-0.6[M/H]+0.3), while larger planets of the sample orbit only stars with [M/H]. These results indicate that while small planets can form in different environments, larger planets preferentially form in metal-rich protoplanetary disks. Exoplanets with P4.3 days orbit on average more metal-rich stars than planets with longer periods. This threshold is smaller than that found for FGK stars (8–10 days) and might be related to M dwarfs having a smaller dust sublimation radius. The distribution of exoplanets with R4R⊕ shows a concentration at orbital periods between 2 and 5 days, which may result from inward orbital migration. There is also a different behavior between single-detected exoplanets and planets from multiplanetary systems, with the latter being found on average around more metal-poor stars, and with planetary radii roughly up to 3 R⊕.
@@ 1 Introduction
Although metallicities have been derived for large numbers of planet-hosting FGK stars, the number of accurate [M/H] determinations for the cooler and lower-mass M-dwarf planet hosts is much smaller. Even though the number of exoplanets detected around M dwarfs to date is considerably smaller than for the FGK stars, this number has been growing, particularly as a result of the TESS mission. Due to both their sheer numbers in the Galaxy and their importance as likely future planet host stars that will be targeted by such missions as the continuing JWST mission or future missions, such as, the Habitable Worlds Observatory (HWO) and PLATO. It becomes increasingly important to provide accurate determinations of the properties of M dwarf stars, especially their metallicities (Rojas-Ayala et al., 2012; Önehag et al., 2012; Newton et al., 2014; Maldonado et al., 2015; Terrien et al., 2015; Lindgren et al., 2016; Lindgren & Heiter, 2017; Souto et al., 2020, 2021; Rajpurohit et al., 2018; Passegger et al., 2022; Birky et al., 2020; Sarmento et al., 2021; Marfil et al., 2021; Iyer et al., 2023; Bello-García et al., 2023; Olander et al., 2025; Behmard et al., 2025). However, there are significant challenges in determining stellar parameters and metallicities for cool stars and fewer works have investigated the metallicities for M dwarf hosts (see Schlaufman & Laughlin 2010; Mann et al. 2013b; Neves et al. 2013; Muirhead et al. 2014; Gaidos & Mann 2014; Souto et al. 2017, 2018; Petigura et al. 2018; Hirano et al. 2018; Dressing et al. 2019; Wang et al. 2024; Hejazi et al. 2024; Melo et al. 2024; Gore et al. 2024; Antoniadis-Karnavas et al. 2024).
@@
A large number of M dwarfs have been observed by the high-resolution near-IR SDSS APOGEE survey (Blanton et al., 2017; Majewski et al., 2017) that has in its DR17 parameters and metallicities for tens of thousands of M dwarfs. However, since APOGEE surveyed primarily red giants in the Galaxy, the APOGEE Stellar Parameters and Abundances Pipeline (ASPCAP; García Pérez et al. 2016) has been tailored to analyze red giants and is not optimized for the analysis of the spectra of M dwarfs, which are much more complex. Previous detailed analyses of APOGEE spectra used dwarf M members of open clusters as benchmarks to verify the accuracy of ASPCAP-derived metallicities, given that (apart from diffusion effects) members of an open cluster are expected to have homogeneous chemical abundances. Souto et al. (2021) analyzed a sample of F, G, K, and M dwarf star members of the open cluster Coma Berenices and showed that the coolest M dwarfs have ASPCAP metallicities that are lower by roughly 0.4 dex when compared to the metallicities of the warmer FGK star cluster members. Such metallicity offsets were later confirmed by Wanderley et al. (2023) who studied Hyades M dwarfs covering a range in effective temperature from 3200 – 3900 K, and finding a median metallicity for the cluster of 0.09 0.03 dex, which is in good agreement with literature results from the optical for solar-type and red-giant stars. ASPCAP metallicity results for the same M dwarf stars showed a trend as a function of the effective temperature of the star. Cooler stars with T3200 K had systematically lower metallicities by up 0.3, which is around 0.4 dex smaller than the median metallicity that Wanderley et al. (2023) obtained for the cluster. The systematic abundance offsets for M dwarf members of open clusters indicated that there are unresolved issues with the APOGEE ASPCAP metallicities, in particular for cool M dwarfs. Alternatively, the spectroscopic analysis methodology from the studies of Souto et al. has been well vetted to produce accurate metallicities for M dwarf stars.
@@
Wanderley et al. (2025) studied a sample of exoplanets having M-dwarf stellar hosts observed by APOGEE, focusing on the discussion of small (R4 R⊕) having orbital periods of P100 days, to probe the exoplanet radius gap, Rgap (Fulton et al. 2017) around cool, low-mass stars. This study detected a well-defined radius gap among the M-dwarf exoplanets at R1.80.2R⊕, a value that was found to remain roughly constant over the orbital period range of P1–20 days; this behavior of Rgap being independent of Porb is in contrast to that found for the hotter and more massive F, G, and K dwarf planet hosts, where the value of Rgap decreases with Porb at a rate of d(logRgap)/d(logPorb) -0.09 to -0.11 (Van Eylen et al., 2018; Martinez et al., 2019). Since the radius gap, which lies between the super-Earth and sub-Neptune exoplanetary regimes, is likely sculpted by combinations of photoevaporation (Lopez & Rice, 2018) / planetary core-powered mass loss (Gupta & Schlichting, 2019), with orbital migration (see also Gaidos et al. 2024; Luque & Pallé 2022). The differing behaviors of Rgap as a function of Porb suggest differences in certain aspects of planetary formation and/or evolution around M dwarfs relative to more massive FGK dwarf stars.
@@
The aim of this paper is to further probe the properties of M dwarf stars by exploring connections with exoplanet properties, such as exoplanet radius, orbital period, and system multiplicity. The first part of this paper investigates relationships of planet properties with host star metallicities and oxygen abundances, which were derived for a subsample of 45 M dwarfs, hosting 65 exoplanets, via quantitative spectroscopic analyses of APOGEE spectra. The second part of this paper analyzes a sample of 246 exoplanets that orbit 188 M dwarf stars (most of this sample was studied for the radius gap of small planets in Wanderley et al. (2025). This paper is organized as follows. In Section 2, we present the sample of exoplanets and their host stars, along with the methodology for deriving stellar metallicities and oxygen abundances, and comparisons of these parameters with results from the literature. In Section 3, we discuss the results, which include the obtained stellar abundances, trends of exoplanetary radii and orbital period with host star metallicities, and trends of exoplanet radii with orbital periods focusing on large planets. Finally, Section 4 summarizes the conclusions.
@@ 2.1.1 Comparisons with the Literature
One of the main goals of this study is to investigate possible trends of the metallicities of M dwarf planet hosts and their exoplanets and, to do so, we need a reliable metallicity scale for the M dwarfs. A comparison between our metallicities with the ASPCAP pipeline (García Pérez et al., 2016) DR17 metallicities is shown in the upper panel of Figure 3. There is a clear trend of the metallicity difference with Teff, with [M/H] ”This Work ASPCAP” reaching 0.3 dex for the coolest stars in our sample, while the offset in metallicity is small for the hottest M dwarf stars with effective temperatures 3900 K. Systematic offsets in the ASPCAP metallicities of cool M dwarfs are well known by the APOGEE ASPCAP team, given that the ASPCAP pipeline was optimized for the analysis of red-giant spectra. As mentioned in the Introduction, metallicity discrepancies have been reported in the literature (Wanderley et al., 2023; Souto et al., 2021) for benchmark M dwarf stars from open clusters.
Benchmark Stars: To further test our metallicity scale, in this study, we also derived stellar parameters and metallicities for 10 benchmark M dwarfs previously studied in Souto et al. (2020, 2022), which are binary stars that have FGK companions. These results are presented in the lower panel of Figure 3. We show the difference, as a function of Teff, between our metallicities for the M dwarfs and the metallicities of the warmer companions that were collected from other works in the literature (Adibekyan et al. 2012; Bensby et al. 2014; Carretta et al. 2013; De Silva et al. 2015; Ghezzi et al. 2010; Lambert & Reddy 2004; Mann et al. 2013a; Mishenina et al. 2008; Ramírez et al. 2007, 2012; Reddy et al. 2006). We find for this comparison a median offset of between the M dwarf and the FGK star metallicities, which is a small difference and generally in line with the assumption of chemical homogeneity of stars born in the same molecular cloud. Different from what is observed in the upper panel of Figure 3, the metallicity differences when using as comparisons results from high-resolution optical studies for warmer stars (bottom panel) do not show a clear trend with the effective temperature, suggesting that our metallicity scale are reliable.
@@ 2.2 Stellar and Planetary Radius Determinations
We used the M versus stellar radius (R) relation in Equation 1 derived in Wanderley et al. (2025) to determine stellar radii for an additional 28 M dwarf stars that were not analyzed in their work.
@@ 3.2 Trends of Exoplanetary Radii and Orbital Period with Metallicity
Figure 5 shows the exoplanetary radii as a function of metallicity (left panel), as well as [O/H] (right panel) for the studied sample of M-dwarf host stars. The 68 planets orbiting M dwarfs displayed in Figure 5 have been identified as being in single (blue filled circles) or multis (red circles), with the largest known exoplanets in multi-planetary systems shown as filled red circles, while the remaining exoplanets from the multi-systems are shown as red open circles. The behavior of Rp with host star metallicity is similar for both the metallicity [M/H] and the oxygen abundance [O/H], both showing an overall positive trend of increasing Rp with [M/H] and [O/H], with the two highest metallicity M dwarfs hosting the two largest planets in our sample: Kepler-45b (hot Jupiter, Rp = 12.6 R⊕), host star metallicity of [M/H]=+0.25; and Kepler-1624 (sub-Saturn, Rp=7.2 R⊕), host star metallicity of [M/H]=+0.26. The vertical dashed line marks the solar metallicity, and we can see that the metal-poor M dwarfs ([M/H]0) host only smaller planets, with Rp less than approximately 2.5 – 3 R⊕, while the metal-rich M dwarfs ([M/H]0) host planets that span the entire range in planetary radii of our sample, with giant planets orbiting more metal-rich hosts. (This is also seen in the Hirano et al. (2018) literature compilation of heterogeneous metallicities). The signature found here for M dwarfs is generally in line with what has been previously found for F, G, and K hosts which showed that Neptune size and smaller exoplanets orbit stars with a range of metallicities, while larger exoplanets tended to be found preferentially around above solar metallicity stars (Buchhave et al., 2012; Sousa et al., 2018; Petigura et al., 2018; Ghezzi et al., 2010; Adibekyan, 2019; Ghezzi et al., 2021). The general increase in Rp with [M/H] (or [O/H]) is found for both single planets, as well as the largest planets in multi-systems. For exoplanets that are the largest members of multiple systems, a quadratic relation of the form: Rp= 2.83 + 4.20 [M/H] + 2.42 [M/H]2 (in R⊕ units, shown by the red dashed line in Figure 5), quantifies the positive correlation between planetary radius and metallicity of the host star. A similar quadratic fit using oxygen abundances is also shown in the right panel for Rp versus [O/H]. The largest single planets for a given metallicity (upper envelope of the blue circles distribution) seem to also follow the same relation as that defined by the largest planets in multis. Notably, there are no planets lying significantly above the red dashed curves and in particular, there is an absence of exoplanets with Rp 2.5 R⊕ in our sample around stars with [M/H] 0. This may indicate a threshold in the maximum planet size with metallicity. Such behavior is also generally seen for the F, G, K hosts (see Figure 3 in Ghezzi et al. 2021), although the threshold in that case was taken as 4.4 R⊕.
@@
A feature of Figure 5 is the different metallicity distributions between the single and multi-systems: of the 33 single system planets, 15 orbit M-dwarf hosts with [M/H]0.0 and 18 have hosts with [M/H]0.0, while for the 12 multi-system planets, the respective numbers are 9 and 3. The statistics remain the same if oxygen is used as the metallicity indicator, with the breakdown for single systems being 17 ([M/H]0.0) and 16 ([M/H]0.0), and 9 and 3, respectively, for multis. This difference in metallicity between single and multi-system M-dwarf hosts can be quantified by mean differences, with the meanSTD stellar metallicity for multi-planet and single-planet systems being, respectively, and , indicating an offset in metallicity between the two populations. A cumulative metallicity distribution between singles and multis is shown in the bottom panels of Figure 5). It is clear that the metallicity distribution of singles versus multis is quite different. Using a two-sample Kolmogorov-Smirnov (K-S) test, we obtain a p-value of 0.012, which indicates at a more than 95 confidence level that the metallicity distributions of multi and single samples are likely not drawn from the same population.
@@
Metallicity differences between single- and multi-planet systems around M dwarfs have also been pointed out by Rodríguez Martínez et al. (2023), using heterogeneous metallicity values from the NASA Exoplanet Archive. Such metallicity effect for M-dwarf hosts is in contrast to the lack of differences in metallicity between single and multi-systems for the hotter (T 4600 - 6500 K) and more massive F, G and K planet hosts from Romero-Mirza & Kempton (2018); Weiss et al. (2018); Ghezzi et al. (2021). In comparisons between singles and multis, it should be pointed out that near-coplanarity is a requirement for the detection of multiple planets via transits, so the differences in [M/H] distributions between multi and single systems could be that a significant fraction of the single systems may be, in fact, multi-systems that are simply less coplanar than the observed metal-poor multi-systems. A reason for such differences in planetary orbital inclinations could be that metal-rich proto-planetary disks enhance the occurrence rates of distant large planets, which can then perturb the orbital planes of more closely orbiting planets; such a scenario is suggested by Pan et al. (2025) and is shown in their figure 4 (right panel). Why such a metallicity-dependent process would lead to contrasting behavior between single- and multi-planet systems around F, G, K dwarfs when compared to the M dwarfs is not known, but suggests differences in the formation and/or evolution of planetary systems around different planet-host stellar masses and luminosities.
@@
We now investigate if there is a correlation between the metallicities of the M dwarf planet hosts and the exoplanet orbital periods, as was previously done for the F, G and K hosts. In the top panel of Figure 6 we show the p-values from the Kolmogorov-Smirnov tests versus the exoplanet orbital period for our sample. To compute the p-values, we separated our sample considering different thresholds in orbital periods in a 0.1-day step. For each orbital period threshold, we divided the sample into a ”short-period” sample and ”long-period” sample, and we performed a K-S test between the two corresponding metallicity distributions, computing its p-value. The lowest p-value obtained was 6.410-5, corresponding to an orbital period of P4.3 days (dashed vertical line in Figure 6). This p-value confirms with a 99 confidence level that the metallicity distributions for exoplanets with Porb longer and shorter than 4.3 days are statistically different.
@@
The bottom panel of Figure 6 also shows the Kernel regression of the mean metallicity (grey curve) as a function of orbital period for our sample. This relation was derived using Equations 2 and 3 from Wilson et al. (2018), adopting a bandwidth of =0.50. The dashed horizontal lines in this figure are the median host star metallicities on each side of the threshold orbital period, with a median value above solar ([M/H] +) for Porb 4.3 days and significantly below solar ([M/H] -0.24) for Porb 4.3 days. Both the Kernel regression and our median metallicities confirm that exoplanets with P4.3 days preferentially orbit more metal-rich stars if compared to exoplanets with larger orbital periods.
@@
Differences in threshold orbital periods between the F, G, and K dwarfs and M dwarfs of, respectively, 8–10 days versus 4.3 days can be mapped into approximate differences in orbital sizes using typical masses for FGK dwarfs and M dwarfs. Using a single mass of M1.0M⊙ for a G dwarf as an example,and a threshold orbital period of P 9 days implies a semi-major axis, a0.08 AU, while for the M dwarfs, a rough mass of M0.4M⊙ and a P4.3 days leads to a0.04 AU – about a factor of two difference in the semi-major axes. One possibility to explain the threshold orbital period as a function of [M/H] that was explored, for example, by Wilson et al. (2018), is that this period occurs at a distance below which dust cannot form (a ‘dust sublimation” radius, Rsub), which is a function of metallicity in the proto-planetary disk (e.g., Millan-Gabet & Monnier 2002). Such a dust sublimation radius for M dwarfs is expected to fall at shorter orbital periods than for FGK stars, because R(Stellar Luminosity)1/2 (Millan-Gabet & Monnier, 2002). The dust sublimation radius hypothesis was discussed in Wilson et al. (2018) but was discarded in their study of FGK hosts because they did not find a significant relation between the metallicity and the effective temperature of the host star in their sample. The lower orbital period for the M dwarfs obtained here in comparison with that for the FGK dwarfs, however, may be a signature of dust sublimation radius differences that appear due to the considerably larger contrast between the luminosities of the M dwarfs relative to the FGK dwarfs, than comparisons within the F, G, and K dwarfs themselves.

### v3

@@ Abstract
One important property in studying the exoplanet population is the host star metallicity ([M/H]). In this study, we derived stellar metallicities and oxygen abundances for 48 M dwarf stars using the near-infrared high-resolution spectra from the SDSS APOGEE survey and synthetic spectra computed in LTE. We also derived and investigated the exoplanetary radii distribution for a larger sample of 246 exoplanets orbiting 188 M dwarf stars. The [M/H] versus [O/M] distribution obtained indicates that our sample is composed mainly of thin disk stars, which follow the behavior of the low-alpha sequence in the Milky Way thin disk. Small planets with radii smaller than 3R⊕ were found around stars with a range of metallicities (-0.6[M/H]+0.3), while larger planets of the sample orbit only stars with [M/H]. These results indicate that while small planets can form in different environments, larger planets preferentially form in metal-rich protoplanetary disks. Exoplanets with P4.3 days orbit on average more metal-rich stars than planets with longer periods. This threshold is smaller than that found for FGK stars (8–10 days) and might be related to M dwarfs having a smaller dust sublimation radius. The distribution of exoplanets with R4R⊕ shows a concentration at orbital periods between 2 and 5 days, which may result from inward orbital migration. There is also a different behavior between single-detected exoplanets and planets from multiplanetary systems, with the latter being found on average around more metal-poor stars, and with planetary radii roughly up to 3 R⊕.
@@ 1 Introduction
Although metallicities have been derived for large numbers of planet-hosting FGK stars, the number of accurate [M/H] determinations for the cooler and lower-mass M-dwarf planet hosts is much smaller. Even though the number of exoplanets detected around M dwarfs to date is considerably smaller than for the FGK stars, this number has been growing, particularly as a result of the TESS mission. Due to both their sheer numbers in the Galaxy and their importance as likely future planet host stars that will be targeted by such missions as the continuing JWST mission or future missions, such as, the Habitable Worlds Observatory (HWO) and PLATO. It becomes increasingly important to provide accurate determinations of the properties of M dwarf stars, especially their metallicities (Rojas-Ayala et al., 2012; Önehag et al., 2012; Newton et al., 2014; Maldonado et al., 2015; Terrien et al., 2015; Lindgren et al., 2016; Lindgren & Heiter, 2017; Souto et al., 2020, 2021; Rajpurohit et al., 2018; Passegger et al., 2022; Birky et al., 2020; Sarmento et al., 2021; Marfil et al., 2021; Iyer et al., 2023; Bello-Garc´ıa et al., 2023; Olander et al., 2025; Behmard et al., 2025). However, there are significant challenges in determining stellar parameters and metallicities for cool stars and fewer works have investigated the metallicities for M dwarf hosts (see Schlaufman & Laughlin 2010; Mann et al. 2013b; Neves et al. 2013; Muirhead et al. 2014; Gaidos & Mann 2014; Souto et al. 2017, 2018; Petigura et al. 2018; Hirano et al. 2018; Dressing et al. 2019; Wang et al. 2024; Hejazi et al. 2024; Melo et al. 2024; Gore et al. 2024; Antoniadis-Karnavas et al. 2024).
@@
A large number of M dwarfs have been observed by the high-resolution near-IR SDSS APOGEE survey (Blanton et al., 2017; Majewski et al., 2017) that has in its DR17 parameters and metallicities for tens of thousands of M dwarfs. However, since APOGEE surveyed primarily red giants in the Galaxy, the APOGEE Stellar Parameters and Abundances Pipeline (ASPCAP; Garc´ıa Pérez et al. 2016) has been tailored to analyze red giants and is not optimized for the analysis of the spectra of M dwarfs, which are much more complex. Previous detailed analyses of APOGEE spectra used dwarf M members of open clusters as benchmarks to verify the accuracy of ASPCAP-derived metallicities, given that (apart from diffusion effects) members of an open cluster are expected to have homogeneous chemical abundances. Souto et al. (2021) analyzed a sample of F, G, K, and M dwarf star members of the open cluster Coma Berenices and showed that the coolest M dwarfs have ASPCAP metallicities that are lower by roughly 0.4 dex when compared to the metallicities of the warmer FGK star cluster members. Such metallicity offsets were later confirmed by Wanderley et al. (2023) who studied Hyades M dwarfs covering a range in effective temperature from 3200 – 3900 K, and finding a median metallicity for the cluster of 0.09 0.03 dex, which is in good agreement with literature results from the optical for solar-type and red-giant stars. ASPCAP metallicity results for the same M dwarf stars showed a trend as a function of the effective temperature of the star. Cooler stars with T3200 K had systematically lower metallicities by up 0.3, which is around 0.4 dex smaller than the median metallicity that Wanderley et al. (2023) obtained for the cluster. The systematic abundance offsets for M dwarf members of open clusters indicated that there are unresolved issues with the APOGEE ASPCAP metallicities, in particular for cool M dwarfs. Alternatively, the spectroscopic analysis methodology from the studies of Souto et al. has been well vetted to produce accurate metallicities for M dwarf stars.
@@
Wanderley et al. (2025) studied a sample of exoplanets having M-dwarf stellar hosts observed by APOGEE, focusing on the discussion of small (R4 R⊕) having orbital periods of P100 days, to probe the exoplanet radius gap, Rgap (Fulton et al. 2017) around cool, low-mass stars. This study detected a well-defined radius gap among the M-dwarf exoplanets at R1.80.2R⊕, a value that was found to remain roughly constant over the orbital period range of P1–20 days; this behavior of Rgap being independent of Porb is in contrast to that found for the hotter and more massive F, G, and K dwarf planet hosts, where the value of Rgap decreases with Porb at a rate of d(logRgap)/d(logPorb) -0.09 to -0.11 (Van Eylen et al., 2018; Martinez et al., 2019). The radius gap for exoplanets around M dwarfs, lies between the super-Earth and sub-Neptune exoplanetary regimes, and is likely sculpted by combinations of photoevaporation (Lopez & Rice, 2018) / planetary core-powered mass loss (Gupta & Schlichting, 2019), with orbital migration (see also Gaidos et al. 2024; Luque & Pallé 2022). The differing behaviors of Rgap as a function of Porb suggest differences in certain aspects of planetary formation and/or evolution around M dwarfs relative to more massive FGK dwarf stars.
@@
The aim of this paper is to further probe the properties of M dwarf stars by exploring connections with exoplanet properties, such as exoplanet radius, orbital period, and system multiplicity. The first part of this paper investigates relationships of planet properties with host star metallicities and oxygen abundances, which were derived for a subsample of 48 M dwarfs, via quantitative spectroscopic analyses of APOGEE spectra. The second part of this paper analyzes a sample of 246 exoplanets that orbit 188 M dwarf stars (most of this sample was studied for the radius gap of small planets in Wanderley et al. (2025). This paper is organized as follows. In Section 2, we present the sample of exoplanets and their host stars, along with the methodology for deriving stellar metallicities and oxygen abundances, and comparisons of these parameters with results from the literature. In Section 3, we discuss the results, which include the obtained stellar abundances, trends of exoplanetary radii and orbital period with host star metallicities, and trends of exoplanet radii with orbital periods focusing on large planets. Finally, Section 4 summarizes the conclusions.
@@ 2.1.1 Comparisons with the Literature
One of the main goals of this study is to investigate possible trends of the metallicities of M dwarf planet hosts and their exoplanets and, to do so, we need a reliable metallicity scale for the M dwarfs. A comparison between our metallicities with the ASPCAP pipeline (Garc´ıa Pérez et al., 2016) DR17 metallicities is shown in the upper panel of Figure 3. There is a clear trend of the metallicity difference with Teff, with [M/H] “This Work ASPCAP” reaching 0.3 dex for the coolest stars in our sample, while the offset in metallicity is small for the hottest M dwarf stars with effective temperatures 3900 K. Systematic offsets in the ASPCAP metallicities of cool M dwarfs are well known by the APOGEE ASPCAP team, given that the ASPCAP pipeline was optimized for the analysis of red-giant spectra. As mentioned in the Introduction, metallicity discrepancies have been reported in the literature (Wanderley et al., 2023; Souto et al., 2021) for benchmark M dwarf stars from open clusters.
Benchmark Stars: To further test our metallicity scale, in this study, we also derived stellar parameters and metallicities for 10 benchmark M dwarfs previously studied in Souto et al. (2020, 2022), which are binary stars that have FGK companions. These results are presented in the lower panel of Figure 3. We show the difference, as a function of Teff, between our metallicities for the M dwarfs and the metallicities of the warmer companions that were collected from other works in the literature (Adibekyan et al. 2012; Bensby et al. 2014; Carretta et al. 2013; De Silva et al. 2015; Ghezzi et al. 2010; Lambert & Reddy 2004; Mann et al. 2013a; Mishenina et al. 2008; Ram´ırez et al. 2007, 2012; Reddy et al. 2006). We find for this comparison a median offset of between the M dwarf and the FGK star metallicities, which is a small difference and generally in line with the assumption of chemical homogeneity of stars born in the same molecular cloud. Different from what is observed in the upper panel of Figure 3, the metallicity differences when using as comparisons results from high-resolution optical studies for warmer stars (bottom panel) do not show a clear trend with the effective temperature, suggesting that our metallicity scale are reliable.
@@ 2.2 Stellar and Planetary Radius Determinations
We used the M versus stellar radius (R) relation in Equation 1 derived in Wanderley et al. (2025) to determine stellar radii for an additional 24 M dwarf stars that were not analyzed in their work.
@@ 3.2 Trends of Exoplanetary Radii and Orbital Period with Metallicity
Figure 5 shows the exoplanetary radii as a function of metallicity (left panel), as well as [O/H] (right panel) for the studied sample of M-dwarf host stars. The 65 planets orbiting 45 M dwarfs displayed in Figure 5 have been identified as being in single (blue filled circles) or multis (red circles), with the largest known exoplanets in multi-planetary systems shown as filled red circles, while the remaining exoplanets from the multi-systems are shown as red open circles. The behavior of Rp with host star metallicity is similar for both the metallicity [M/H] and the oxygen abundance [O/H], both showing an overall positive trend of increasing Rp with [M/H] and [O/H], with the two highest metallicity M dwarfs hosting the two largest planets in our sample: Kepler-45b (hot Jupiter, Rp = 12.6 R⊕), host star metallicity of [M/H]=+0.25; and Kepler-1624 b (sub-Saturn, Rp=7.2 R⊕), host star metallicity of [M/H]=+0.26. The vertical dashed line marks the solar metallicity, and we can see that the metal-poor M dwarfs ([M/H]0) host only smaller planets, with Rp less than approximately 2.5 – 3 R⊕, while the metal-rich M dwarfs ([M/H]0) host planets that span the entire range in planetary radii of our sample, with giant planets orbiting more metal-rich hosts. (This is also seen in the Hirano et al. (2018) literature compilation of heterogeneous metallicities). The signature found here for M dwarfs is generally in line with what has been previously found for F, G, and K hosts which showed that Neptune size and smaller exoplanets orbit stars with a range of metallicities, while larger exoplanets tended to be found preferentially around stars with [M/H]0 (Buchhave et al., 2012; Sousa et al., 2018; Petigura et al., 2018; Ghezzi et al., 2010; Adibekyan, 2019; Ghezzi et al., 2021). The general increase in Rp with [M/H] (or [O/H]) is found for both single planets, as well as the largest planets in multi-systems. For exoplanets that are the largest members of multiple systems, a quadratic relation of the form: Rp= 2.83 + 4.20 [M/H] + 2.42 [M/H]2 (in R⊕ units, shown by the red dashed line in Figure 5), quantifies the positive correlation between planetary radius and metallicity of the host star. A similar quadratic fit using oxygen abundances is also shown in the right panel for Rp versus [O/H]. The largest single planets for a given metallicity (upper envelope of the blue circles distribution) seem to also follow the same relation as that defined by the largest planets in multis. Notably, there are no planets lying significantly above the red dashed curves and in particular, there is an absence of exoplanets with Rp 2.5 R⊕ in our sample around stars with [M/H] 0. This may indicate a threshold in the maximum planet size with metallicity. Such behavior is also generally seen for the F, G, K hosts (see Figure 3 in Ghezzi et al. 2021), although the threshold in that case was taken as 4.4 R⊕.
@@
A feature of Figure 5 is the different metallicity distributions between the single and multi-systems: of the 32 single system planets, 14 orbit M-dwarf hosts with [M/H]0.0 and 18 have hosts with [M/H]0.0, while for the 13 multi-systems, the respective numbers are 10 and 3. The statistics remain similar if oxygen is used as the metallicity indicator, with the breakdown for single systems being 16 ([O/H]0.0) and 16 ([O/H]0.0), and 11 and 2, respectively, for multis. This difference in metallicity between single and multi-system M-dwarf hosts can be quantified by mean differences, with the meanSTD stellar metallicity for multi-planet and single-planet systems being, respectively, and , indicating an offset in metallicity between the two populations. A cumulative metallicity distribution between singles and multis is shown in the bottom panels of Figure 5). It is clear that the metallicity distribution of singles versus multis is quite different. Using a two-sample Kolmogorov-Smirnov (K-S) test, we obtain a p-value of 0.002, which indicates at a 99 confidence level that the metallicity distributions of multi and single samples are likely not drawn from the same population.
@@
Metallicity differences between single- and multi-planet systems around M dwarfs have also been pointed out by Rodr´ıguez Mart´ınez et al. (2023), using heterogeneous metallicity values from the NASA Exoplanet Archive. Such metallicity effect for M-dwarf hosts is in contrast to the lack of differences in metallicity between single and multi-systems for the hotter (T 4600 - 6500 K) and more massive F, G and K planet hosts from Romero-Mirza & Kempton (2018); Weiss et al. (2018); Ghezzi et al. (2021). In comparisons between singles and multis, it should be pointed out that near-coplanarity is a requirement for the detection of multiple planets via transits, so the differences in [M/H] distributions between multi and single systems could be that a significant fraction of the single systems may be, in fact, multi-systems that are simply less coplanar than the observed metal-poor multi-systems. A reason for such differences in planetary orbital inclinations could be that metal-rich proto-planetary disks enhance the occurrence rates of distant large planets, which can then perturb the orbital planes of more closely orbiting planets; such a scenario is suggested by Pan et al. (2025) and is shown in their figure 4 (right panel). Why such a metallicity-dependent process would lead to contrasting behavior between single- and multi-planet systems around F, G, K dwarfs when compared to the M dwarfs is not known, but it suggests differences in the formation and/or evolution of planetary systems around different planet-host stellar masses and luminosities.
@@
We now investigate if there is a correlation between the metallicities of the M dwarf planet hosts and the exoplanet orbital periods, as was previously done for the F, G and K hosts. In the top panel of Figure 6 we show the p-values from the Kolmogorov-Smirnov tests versus the exoplanet orbital period for our sample. To compute the p-values, we separated our sample considering different thresholds in orbital periods in a 0.1-day step. For each orbital period threshold, we divided the sample into a “short-period” sample and “long-period” sample, and we performed a K-S test between the two corresponding metallicity distributions, computing its p-value. The lowest p-value obtained was 6.410-5, corresponding to an orbital period of P4.3 days (dashed vertical line in Figure 6). This p-value confirms with a 99 confidence level that the metallicity distributions for exoplanets with Porb longer and shorter than 4.3 days are statistically different.
@@
The bottom panel of Figure 6 also shows the Kernel regression of the mean metallicity (gray curve) as a function of orbital period for our sample. This relation was derived using Equations 2 and 3 from Wilson et al. (2018), adopting a bandwidth of =0.50. The dashed horizontal lines in this figure are the median host star metallicities on each side of the threshold orbital period, with a median value above solar ([M/H] +) for Porb 4.3 days and significantly below solar ([M/H] -0.24) for Porb 4.3 days. Both the Kernel regression and our median metallicities confirm that exoplanets with P4.3 days preferentially orbit more metal-rich stars if compared to exoplanets with larger orbital periods.
@@
Differences in threshold orbital periods between the F, G, and K dwarfs and M dwarfs of, respectively, 8–10 days versus 4.3 days can be mapped into approximate differences in orbital sizes using typical masses for FGK dwarfs and M dwarfs. Using a single mass of M1.0M⊙ for a G dwarf as an example,and a threshold orbital period of P 9 days implies a semi-major axis, a0.08 AU, while for the M dwarfs, a rough mass of M0.4M⊙ and a P4.3 days leads to a0.04 AU – about a factor of two difference in the semi-major axes. One possibility to explain the threshold orbital period as a function of [M/H] that was explored, for example, by Wilson et al. (2018), is that this period occurs at a distance below which dust cannot form (a “dust sublimation” radius, Rsub), which is a function of metallicity in the proto-planetary disk (e.g., Millan-Gabet & Monnier 2002). Such a dust sublimation radius for M dwarfs is expected to fall at shorter orbital periods than for FGK stars, because R(Stellar Luminosity)1/2 (Millan-Gabet & Monnier, 2002). The dust sublimation radius hypothesis was discussed in Wilson et al. (2018) but was discarded in their study of FGK hosts because they did not find a significant relation between the metallicity and the effective temperature of the host star in their sample. The lower orbital period for the M dwarfs obtained here in comparison with that for the FGK dwarfs, however, may be a signature of dust sublimation radius differences that appear due to the considerably larger contrast between the luminosities of the M dwarfs relative to the FGK dwarfs, than comparisons within the F, G, and K dwarfs themselves.